Alexander Miles

amiles@email.arizona.edu

Objective

To pursue education and research opportunities in the field of optical materials, specifically glasses and ceramics for high-performance optics, solar energy, and remote sensing applications.

Education

University of Arizona, AZ  Major GPA 3.954/4.0, Overall 3.884/4.0  Graduation - May 2011

Pursuing a Bachelor of Science, Materials Science Engineering Pursuing a Bachelor of Science, Optical Sciences and Engineering

Professional Experience

SCHOTT AG - Advanced Optics Division - Mainz, Germany Summer 2011 Praktikant (Intern) to Dr. Reiter and Dr. Schreder

• Designed battery of test melts and experiments to diagnose negative surface reactions.

• Performed and reviewed German-English translations for technical accuracy and flow.

• Wrote and ran simulations of thermal processing of specific glass types.

Lunar and Planetary Sciences / NASA Space Grant Academic Year 2009-2010 Research Assistant and Programmer for Dr. Robert Erdmann

• Deconstructed and rebuilt exiting modeling code to simulate new experiments.

• Visualized large data sets for further analysis.

University of Arizona Materials Science Department Summer 2009 Student Research Assistant to Dr. Barrett G. Potter, Jr.

• Run batches of time intensive experiments focused on investigating the effect of various parameters on stress-linked optical phenomenon in KDP crystals.

• Undertook extensive documentation and characterization of experiments and samples.

• Revised existing sample preparation procedure, reducing scatter due to inconsistent surface finish.

University of Arizona Materials Science Department Summer 2008 Student Research Assistant to Dr. Pierre Lucas

• Designed and ran experiments to determine ideal parameters for induced photo-relaxation of calcogenide glasses.

• Used results to explore range of photoplastic phenomenon possible within a subset of glasses.

• Trained on and utilized laser systems, surface characterization, and differential scanning calorimetry.

Steward Observatory Imaging Technology Lab September 2007 - December 2009 Student Research Assistant

• Trained on myriad machines for routine and experimental tasks.

• Ran batteries of tests on optical coatings for surface roughness and suitability for use on CCDs.

• Planned and undertook research project on bending silicon detectors to a focal plane.

• Wrote programs (Python) to remotely manage optical and computer systems.

Technical Skills

LabView, CODE V, SolidWorks, Microsoft Office Suite 2010, Adobe Creative Suite CS4, OpenOffice Suite, MATLAB, PHP, MySQL, HTML, CSS, Javascript, Python, SciPy, NumPy, L^{A}TEX, Drupal 5, Debian Linux, GNU Emacs, . Experienced in chemical and digital photography, glassblowing in both German and Venetian style, web design, and workstation repair.

Memberships and Honors NASA Space Grant Recipient University of Arizona da Vinci Scholar 2010 Thomas G. Chapman Scholarship Nominee SPIE Member, University of Arizona Chapter SCHOTT Advanced Optics Scholarship Recipient Material Advantage, University of Arizona Chapter Dean’s List with Distinction, University of Arizona President of Keramos, University of Arizona Chapter Presidents Award for Excellence, University of Arizona

Programming Lead for SEDS Rocketry (Students for the Exploration and Development of Space)

References and additional employment history available upon request.